OpenStudy (anonymous):

Identify all of the following solutions of square root of x minus 8 end root plus 8 equals x.

OpenStudy (anonymous):

is this the problem: \(\large \sqrt{x-8}+8=x \) ???

OpenStudy (anonymous):

if yes, subtract 8 from both sides, then square both sides: \(\large \sqrt{x-8}+8-8=x-8\) \(\large \sqrt{x-8}=x-8\) \(\large x-8=(x-8)^2\) \(\large x-8=x^2-16x+64\) can you take it from here?